OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 15 and the last term is –4,199,040? I know how to find sum with the equation I have but I can't find common ratio. Can someone please explain how to find it.

OpenStudy (amistre64):

the common ratio can be determined as: an = a1 r^(n-1) , assuming we start at n=1 divide by a1, and root off the n-1 to determine r (n-1)rt (an/a1) = r

OpenStudy (amistre64):

n=8 a1 =15 a8 = –4,199,040 ------------------------- (8-1)rt (a8/a1) = r 7rt ( –4,199,040/15) = r

OpenStudy (anonymous):

Thank you I get it now :)

OpenStudy (amistre64):

:) good luck